OpenStudy (anonymous):

Where does the horizontal asymptote lie for f(x) = 5 − 2x/2 − x? y = 1 y = 2 y = 3 y = 5

zepdrix (zepdrix):

\[\Large\rm f(x)=\frac{-2x+5}{-x+2}\]We use limits to find out what's happening on the `ends` of a function. This tells us about horizontal asymptotes. The idea is basically, let's plug in a huge value for x and see what's happening way to the right side of the graph. If we plug in x=12bajillion we get,\[\Large\rm \approx\frac{-2(12bajillion)}{-(12bajillion)}\]The +5 and +2 are sooooo insignificant at this point that we won't even worry about them.

zepdrix (zepdrix):

If you cancel some stuff out, you'll see that we're approaching the value \(\Large\rm 2\), yes?

zepdrix (zepdrix):

So the shortcut is ... when the leading terms (highest power of x) are equal, the horizontal asymptote will be the `ratio` of their coefficients.\[\Large\rm \lim_{x\to\infty}\frac{-2x+5}{-x+2}=\frac{-2}{-1}\]